OpenStudy (anonymous):

12b^2-23b-2=0 Solve Equation

OpenStudy (anonymous):

b=2 or b=-1/12

OpenStudy (anonymous):

Can u please show how me the steps

OpenStudy (anonymous):

Sure 12b^2-24b+b-2=0 12b(b-2)+1(b-2)=0 (12b+1)(b-2)=0 so either b=0 or b=-1/12

OpenStudy (anonymous):

0h it was wondering why the 23b turn into 24b?

OpenStudy (anonymous):

Can you see why now then....

OpenStudy (anonymous):

no i ment why does the 23b turn into 24b?

OpenStudy (radar):

-23b=-24b+b

OpenStudy (anonymous):

Yeap -23b=-24b+b

OpenStudy (radar):

Perfectly legal, and done in order to factor.

OpenStudy (radar):

Saved you from not using that danged old quadratic formula!

OpenStudy (anonymous):

what about 12t^2+5t-2=0

OpenStudy (radar):

Let me look at it and see if we can escape the "formula"

OpenStudy (anonymous):

12t^2+8t-3t-2=0

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

Now factorise

OpenStudy (radar):

t=-1/2 t=1/3

OpenStudy (anonymous):

either t=1/4 or -2/3

OpenStudy (anonymous):

do you guys have a shortcut for these problems?

OpenStudy (anonymous):

Sorry Radar, your one is incorrect!

OpenStudy (radar):

Oops, I see now those are not the correct factors.

OpenStudy (radar):

(3t+2)(4t-1) are the correct factors (3t+2)=0 t=-2/3 4t-1=0 t=1/4

OpenStudy (radar):

Thanks roshan2004, I got off the rail.

OpenStudy (anonymous):

My pleasure..........